What 2 numbers have a product of 128 and a sum of 48?
I don't believe it has a solution with whole numbers anyway.

Answer by jim_thompson5910(35256)

To find these two numbers, we need to list all of the factors of 128.


Factors of 128:
1,2,4,8,16,32,64,128
-1,-2,-4,-8,-16,-32,-64,-128


Note: list the negative of each factor. This will allow us to find all possible combinations.


These factors pair up and multiply to 128.
1*128 = 128
2*64 = 128
4*32 = 128
8*16 = 128
(-1)*(-128) = 128
(-2)*(-64) = 128
(-4)*(-32) = 128
(-8)*(-16) = 128

Now let's add up each pair of factors to see if one pair adds to the middle coefficient 48:


First NumberSecond NumberSum
11281+128=129
2642+64=66
4324+32=36
8168+16=24
-1-128-1+(-128)=-129
-2-64-2+(-64)=-66
-4-32-4+(-32)=-36
-8-16-8+(-16)=-24



From the table, we can see that there are no pairs of numbers which add to 48.


So you are correct. There are no two rational numbers which multiply to 128 and add to 48.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim